OpenStudy (anonymous):

what is the absolute value of -2 +3i

OpenStudy (anonymous):

absolute value of a+bi is sqrt(a^2 + b^2). As an example, the absolute value of 3 - 4i is sqrt(3^2 + (-4)^2)= sqrt (9+16) = sqrt(25) = 5.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Thank you! :)